### v3

@@ Acknowledgments
    Table A3: Results for different MTO methods and optimizers on CityScapes[9] using SegNet [1]. The best score per metric is highlighted for each MTO method as well as over all methods and optimizers. While different MTO methods perform best over the distinct metrics, models trained with Adam outperform those based on SGD+mom in most direct comparisons. On the overall $\Delta_{m}$-metric, Adam shows superior performance for all MTO methods, in some cases even with a high margin. Best performance for each metric was also achieved by using Adam.
    (c) NYUv2,     SegNet
    (d) NYUv2,          DeepLabV3
@@
    (e) CelebA, ResNet50
    Figure A10: Count of positive/ negative/ zero components in gradients. For each gradient corresponding to one task and one sample, we count the number of positive and negative as well as zero-valued scalar entries. As expected, the amount of positive and negative values is approximately equal during the entire time of training along all experiments. For experiments on CityScapes and CelebA, we found that an increasing number of network components would not receive gradient updates anymore starting at some point in training. This especially holds for CelebA, were we assume to happen because of the ReLU activation functions used in ResNet18.
    (a) CityScapes, SegNet
    (b) CityScapes, DeepLabV3
    (c) NYUv2, SegNet
    (d) NYUv2, DeepLabV3
    (e) CelebA, ResNet50
    Figure A11: Loss over experiments for gradient similarity experiment. We visualize the loss with respect to the training, validation, and test set during the training process of the gradient similarity experiments.
    Appendix A6 Additional results for out-of distribution generalization
    We present additional results on generalization performance to corrupted data on CityScapes with SegNet as well as on NYUv2 with either choice of network in Sec. 4.3 and Secs. 4.3 and 4.3. We further show absolute scores for all setups in Secs. 4.3 and 4.3
    (a) CityScapes, SegNet
    (b) NYUv2, SegNet
    (c) NYUv2, DeepLabV3
@@
    SegNet RLW 0.3744 0.7321 0.0526 96.26
    SegNet IMTL 0.3684 0.7363 0.0479 78.92
    SegNet PCGrad 0.3678 0.7310 0.0515 94.44
    SegNet CAGrad 0.3769 0.7397 0.0479 92.75
@@
    DeepLabV3 RLW 0.5634 0.8557 0.0270 103.22
    DeepLabV3 IMTL 0.5769 0.8617 0.0244 101.26
    DeepLabV3 PCGrad 0.5742 0.8609 0.0263 105.46
    DeepLabV3 CAGrad 0.5690 0.8576 0.0259 104.90
@@
    SegNet RLW 0.2182 0.4587 0.7718 0.2957 37.09 34.39 0.1165 0.3045 0.4298
    SegNet IMTL 0.2190 0.4743 0.7804 0.2885 33.89 29.67 0.1609 0.3812 0.5085
    SegNet PCGrad 0.2279 0.4751 0.7790 0.2900 35.34 31.90 0.1388 0.3423 0.4696
    SegNet CAGrad 0.2324 0.4850 0.7822 0.2897 33.50 29.10 0.1648 0.3908 0.5183
@@
    DeepLabV3 RLW 0.3861 0.6338 0.5290 0.2081 28.33 23.00 0.2566 0.4964 0.6162
    DeepLabV3 IMTL 0.3773 0.6323 0.5208 0.1991 27.27 21.34 0.2830 0.5273 0.6412
    DeepLabV3 PCGrad 0.3932 0.6409 0.5168 0.2025 28.05 22.44 0.2659 0.5063 0.6238
    DeepLabV3 CAGrad 0.3738 0.6289 0.5266 0.2022 27.68 21.86 0.2754 0.5177 0.6324

### v5

@@ Acknowledgments
Table A4: Results for different MTO methods and optimizers on Cityscapes[9] using SegNet [1]. The best score per metric is highlighted for each MTO method as well as over all methods and optimizers. While different MTO methods perform best over the distinct metrics, models trained with Adam outperform those based on SGD+mom in most direct comparisons. On the overall $\Delta_{m}$ metric, Adam shows superior performance for all MTO methods, in some cases even with a high margin. Best performance for each metric was also achieved by using Adam.
Sem.Seg. Depth
MTO Optimizer lr mIoU $\uparrow$ pixAcc $\uparrow$ AbsErr $\downarrow$ RelErr $\downarrow$ DeltaM $\downarrow$
STL adam 0.7122 0.9221 0.0134 29.88
EW adam 0.005 0.6898 0.9165 0.0196 109.84 79.43 $\pm$3.68
EW signSGD 0.001 0.7013 0.9174 0.0210 115.67 86.52 $\pm$6.06
EW sgd 0.1 0.6967 0.9179 0.0216 113.82 86.24 $\pm$1.97
[…]
UW signSGD 0.001 0.6506 0.8997 0.0166 53.61 28.54 $\pm$7.83
UW sgd 0.01 0.6750 0.9110 0.0219 114.67 88.39 $\pm$1.35
[…]
RLW signSGD 0.0005 0.6962 0.9169 0.0204 112.46 82.90 $\pm$5.32
RLW sgd 0.1 0.6918 0.9156 0.0227 113.59 88.16 $\pm$0.67
[…]
IMTL signSGD 0.005 0.6659 0.9070 0.0197 101.90 74.03 $\pm$14.62
IMTL sgd 0.01 0.6716 0.9107 0.0230 114.38 90.21 $\pm$0.74
[…]
PCGrad signSGD 0.001 0.6929 0.9170 0.0210 113.32 84.74 $\pm$3.46
PCGrad sgd 0.1 0.6972 0.9176 0.0235 107.06 84.09 $\pm$0.93
[…]
CAGrad signSGD 0.001 0.6883 0.9138 0.0182 107.06 74.68 $\pm$18.91
CAGrad sgd 0.1 0.6896 0.9156 0.0205 115.52 85.88 $\pm$0.31
[…]
AlignedMTL signSGD 0.001 0.7010 0.9189 0.0180 102.41 69.77 $\pm$24.91
AlignedMTL sgd 0.1 0.6684 0.9103 0.0224 113.11 88.37 $\pm$0.53
[…]
MTL-IO sgd 0.1 0.6957 0.9172 0.0233 110.19 86.46 $\pm$1.63
Table A5: Results for different MTO methods and optimizers on Cityscapes [9] using DeepLabV3+ [5]. The best score per metric is highlighted for each MTO method as well as over all methods and optimizers. Adam is Pareto dominant over SGD+mom in a direct pairwise comparison across all MTO methods. Here, we would also like to highlight that MTL can outperform STL as suggested by [4].
Sem.Seg. Depth
MTO Optimizer lr mIoU $\uparrow$ pixAcc $\uparrow$ AbsErr $\downarrow$ RelErr $\downarrow$ DeltaM $\downarrow$
STL adam 0.7203 0.9253 0.0132 47.37
EW adam 0.001 0.7247 0.9268 0.0128 47.73 -0.80 $\pm$0.69
EW signSGD 0.0005 0.7188 0.9249 0.0134 52.47 3.07 $\pm$2.70
EW sgd 0.05 0.7100 0.9217 0.0174 120.34 46.88 $\pm$2.02
[…]
UW signSGD 0.001 0.7172 0.9244 0.0127 48.31 -0.35 $\pm$2.23
UW sgd 0.005 0.7003 0.9187 0.0171 116.09 44.47 $\pm$1.91
[…]
RLW signSGD 0.001 0.7185 0.9244 0.0133 54.75 4.09 $\pm$0.39
RLW sgd 0.05 0.7070 0.9205 0.0176 123.13 48.84 $\pm$1.79
[…]
IMTL signSGD 0.001 0.7177 0.9243 0.0128 49.61 0.57 $\pm$3.75
IMTL sgd 0.005 0.7027 0.9192 0.0185 122.37 50.33 $\pm$3.30
[…]
PCGrad signSGD 0.001 0.7187 0.9248 0.0132 53.35 3.28 $\pm$1.41
PCGrad sgd 0.05 0.7083 0.9212 0.0173 122.52 47.90 $\pm$3.68
[…]
CAGrad signSGD 0.001 0.7154 0.9243 0.0128 49.06 0.37 $\pm$0.41
CAGrad sgd 0.1 0.7096 0.9220 0.0172 120.24 46.58 $\pm$2.21
[…]
AlignedMTL signsgd 0.001 0.7169 0.9241 0.0127 47.56 -0.70 $\pm$0.61
AlignedMTL sgd 0.1 0.7096 0.9218 0.0174 120.38 46.94 $\pm$1.19
[…]
MTL-IO sgd 0.05 0.7098 0.9217 0.0187 120.61 49.57 $\pm$2.28
Table A6: Results for different MTO methods and optimizers on NYUv2[39] using SegNet [1]. The best score per metric is highlighted for each MTO method as well as over all methods and optimizers. Using Adam yields in superior performance in the majority of cases, both when considering the individual metrics and the overall $\Delta_{m}$ metric. We note that $\Delta_{m}$ is more effected by the normal task due to the higher number of corresponding metrics as can be observed in the case of UW.
Sem.Seg. Depth Normal
[…]
STL adam 0.392 0.646 0.607 0.258 24.74 18.49 0.308 0.582 0.700
EW adam 0.0001 0.398 0.650 0.530 0.212 29.53 25.02 0.216 0.454 0.584 10.08 $\pm$2.84
EW signSGD 0.0001 0.389 0.647 0.543 0.220 30.18 25.93 0.204 0.438 0.569 12.75 $\pm$2.24
EW sgd 0.01 0.384 0.644 0.551 0.227 30.12 25.72 0.212 0.443 0.571 12.84 $\pm$0.65
[…]
UW signSGD 0.0001 0.384 0.634 0.535 0.218 28.36 23.45 0.233 0.482 0.612 8.05 $\pm$1.18
UW sgd 0.05 0.383 0.642 0.551 0.218 27.07 21.66 0.259 0.516 0.644 4.46 $\pm$2.13
[…]
RLW signSGD 0.0001 0.375 0.636 0.544 0.226 29.38 24.77 0.217 0.458 0.589 11.42 $\pm$1.02
RLW sgd 0.05 0.369 0.633 0.572 0.233 31.21 27.17 0.198 0.420 0.546 16.82 $\pm$1.16
[…]
IMTL signSGD 0.0001 0.374 0.639 0.527 0.212 26.26 20.65 0.272 0.537 0.663 1.97 $\pm$0.832
IMTL sgd 0.05 0.396 0.656 0.532 0.215 26.16 20.38 0.277 0.542 0.666 0.72 $\pm$0.68
[…]
PCGrad signSGD 0.0001 0.394 0.649 0.556 0.219 28.98 24.25 0.223 0.467 0.598 9.59 $\pm$1.02
PCGrad sgd 0.01 0.389 0.644 0.547 0.223 29.76 25.22 0.214 0.451 0.580 11.61$\pm$0.36
[…]
CAGrad signSGD 0.0001 0.393 0.648 0.546 0.217 26.21 20.58 0.272 0.538 0.665 1.67 $\pm$0.61
CAGrad sgd 0.05 0.400 0.656 0.547 0.223 26.39 20.78 0.268 0.534 0.662 2.06 $\pm$0.65
[…]
AlignedMTL signSGD 0.0001 0.363 0.637 0.528 0.211 25.86 20.02 0.282 0.550 0.674 0.90 $\pm$0.54
AlignedMTL sgd 0.05 0.371 0.642 0.531 0.214 26.41 20.71 0.272 0.535 0.661 2.33 $\pm$1.00
[…]
MTL-IO sgd 0.01 0.377 0.637 0.551 0.225 30.05 25.50 0.215 0.447 0.574 12.69$\pm$1.21
Table A7: Results for different MTO methods and optimizers on NYUv2[39] using DeepLabV3+ [5]. The best score per metric is highlighted for each MTO method as well as over all methods and optimizers. We note a full dominance of Adam over SGD+mom on both the depth and normal tasks as well as on the $\Delta_{m}$ metric. Overall, best results for all metrics were also achieved using Adam as optimizer.
Sem.Seg. Depth Normal
[…]
STL adam 0.552 0.767 0.365 0.152 21.16 14.52 0.402 0.668 0.768
EW adam 0.0001 0.552 0.770 0.352 0.143 22.54 16.05 0.366 0.634 0.742 2.70 $\pm$0.12
EW signSGD 0.0001 0.554 0.766 0.362 0.147 22.69 16.15 0.364 0.630 0.739 3.60 $\pm$0.38
EW sgd 0.01 0.556 0.769 0.365 0.147 23.35 16.76 0.352 0.617 0.727 5.21 $\pm$0.18
[…]
UW signSGD 0.0005 0.526 0.748 0.366 0.146 22.18 15.46 0.380 0.645 0.749 2.90 $\pm$0.79
UW sgd 0.01 0.557 0.769 0.364 0.148 22.37 15.69 0.374 0.641 0.746 2.57 $\pm$0.56
[…]
RLW signSGD 0.0001 0.557 0.768 0.363 0.146 22.40 15.84 0.371 0.638 0.745 2.74 $\pm$0.30
RLW sgd 0.05 0.541 0.762 0.368 0.151 23.00 16.44 0.357 0.625 0.734 5.15 $\pm$0.54
[…]
IMTL signSGD 0.0001 0.552 0.767 0.359 0.146 21.81 15.22 0.385 0.652 0.756 1.12 $\pm$0.16
IMTL sgd 0.05 0.545 0.762 0.367 0.150 21.63 14.98 0.390 0.658 0.760 1.30 $\pm$0.40
[…]
PCGrad signSGD 0.0001 0.553 0.766 0.360 0.148 22.61 16.11 0.365 0.631 0.739 3.55 $\pm$0.13
PCGrad sgd 0.01 0.556 0.769 0.361 0.147 23.36 16.79 0.351 0.617 0.728 5.10 $\pm$0.28
[…]
CAGrad signSGD 0.0001 0.550 0.766 0.361 0.146 21.94 15.27 0.383 0.650 0.754 1.49 $\pm$0.12
CAGrad sgd 0.05 0.556 0.768 0.360 0.144 21.77 14.99 0.391 0.657 0.759 0.42 $\pm$0.07
[…]
AlignedMTL signSGD 0.0005 0.527 0.751 0.360 0.144 21.35 14.60 0.400 0.666 0.765 0.25 $\pm$0.21
AlignedMTL sgd 0.1 0.550 0.766 0.367 0.150 21.59 14.88 0.393 0.659 0.761 0.95 $\pm$0.62
[…]
MTL-IO sgd 0.01 0.554 0.768 0.362 0.146 23.31 16.75 0.352 0.617 0.728 5.01 $\pm$0.38
Appendix A5 Additional gradient alignment results
We report extended evaluation on gradient simlarity in MTL and STL.
Alternatively to a single sample/ task, we consider the average gradient in App. A4. In Fig. A6, we differentiate between conflicting and supporting gradient pairs when evaluating the cosine similarity. Figure A7 shows the evaluation of the scalar product as an combined measure of similarity in gradient direction and magnitude.
(c) Cityscapes, SegNet
(d) Cityscapes,      DeepLabV3
(e) NYUv2,     SegNet
(f) NYUv2,          DeepLabV3
@@
Appendix A6 Robustness of multi-task representations on corrupted data
In this additonal part of our analysis, we investigate whether features learned for multiple tasks generalize better to corrupted data compared to those learned for single tasks only.
Motivation: In his seminal paper, Caruana gives preliminary evidence that MTL provides stronger features and avoids spurious correlations (referred to better attribute selection[4]. More recently, spurious correlations have often been directly connected with robustness [13, 18]. Results from current literature on the robustness of MTL features are mixed. While MTL is stated to increase the adversarial- and noise-robustness over STL[24, 37, 54], others argue features selected by MTL could be more likely to be non-causal and, therefore, less robust [17, 2]. Here, we further examine whether MTL features lead to better robustness. We would like to nuance that we do not consider the transferability of representations, e.g., to new tasks, but solely focus on the claim that the MTL trained features are more robust w.r.t. different inputs.
@@
RLW 0.3744 0.7321 0.0526 96.26
IMTL 0.3684 0.7363 0.0479 78.92
PCGrad 0.3678 0.7310 0.0515 94.44
CAGrad 0.3769 0.7397 0.0479 92.75
@@
RLW 0.5634 0.8557 0.0270 103.22
IMTL 0.5769 0.8617 0.0244 101.26
PCGrad 0.5742 0.8609 0.0263 105.46
CAGrad 0.5690 0.8576 0.0259 104.90
@@
RLW 0.2182 0.4587 0.7718 0.2957 37.09 34.39 0.1165 0.3045 0.4298
IMTL 0.2190 0.4743 0.7804 0.2885 33.89 29.67 0.1609 0.3812 0.5085
PCGrad 0.2279 0.4751 0.7790 0.2900 35.34 31.90 0.1388 0.3423 0.4696
CAGrad 0.2324 0.4850 0.7822 0.2897 33.50 29.10 0.1648 0.3908 0.5183
@@
RLW 0.3861 0.6338 0.5290 0.2081 28.33 23.00 0.2566 0.4964 0.6162
IMTL 0.3773 0.6323 0.5208 0.1991 27.27 21.34 0.2830 0.5273 0.6412
PCGrad 0.3932 0.6409 0.5168 0.2025 28.05 22.44 0.2659 0.5063 0.6238
CAGrad 0.3738 0.6289 0.5266 0.2022 27.68 21.86 0.2754 0.5177 0.6324
@@
Results: Appendix A4 shows $\delta_{t}$ for different corruption types for all network architectures and datasets with EW averaged over five corruption levels and three seeds. Considering the combination DeepLabV3 and Cityscapes, on the semantic segmentation tasks, the STL models show a slightly lower decrease in performance on the corrupted data than MTL ($\delta_{t}>0$ more often; shaded in red), indicating that the features learned for these respective tasks can better generalize to corrupted data. In contrast, the MTL model shows significant better relative performance on the depth task ($\delta_{t}<0$ more often; shaded green). Comparing these observations to other dataset+network combinations, we find a strong robustness of MTL features for both setups of CityScape+SegNet and NYUv2+DeepLabV3 over all tasks. However, evaluation on NYUv2+SegNet resulted on average in better performance on STL features, especially for the depth task. Furthermore, we see little evidence for general higher robustness against certain types of corruption (e.g., higher robustness against weather conditions) for either MTL or STL across all setups.
The results of other MTO methods (App. A4 and A4) indicate that it depends less on the used method but more on the choice of dataset and network architecture whether some tasks would benefit from MTL for learning more robust features. Averaged absolute scores can be found in App. A4 and A4.